# Investigating the Impact of Project Risks on Employee Turnover Intentions in the IT Industry of Pakistan

### v3

@@ Abstract
Employee turnover remains a pressing issue within high-tech sectors such as IT firms and research centers, where organizational success heavily relies on the skills of their workforce. Intense competition and a scarcity of skilled professionals in the industry contribute to a perpetual demand for highly qualified employees, posing challenges for organizations to retain talent. While numerous studies have explored various factors affecting employee turnover in these industries, their focus often remains on overarching trends rather than specific organizational contexts. In particular, within the software industry, where project-specific risks can significantly impact project success and timely delivery, understanding their influence on job satisfaction and turnover intentions is crucial. This study aims to investigate the influence of project risks in the IT industry on job satisfaction and employee turnover intentions. Furthermore, it examines the role of both external and internal social links in shaping perceptions of job satisfaction.
@@
Project risks, Employee turnover intention, Extreme learning machine, Job satisfaction, IT industry
@@ 1 Introduction
Employee turnover poses a significant challenge in the high-tech industry, where skilled personnel are indispensable for success. This is particularly evident in the IT sector, where employees play a pivotal role throughout various project phases, from requirement gathering to development. The departure of key employees can disrupt projects, leading to delays or even failure. While early perspectives attributed turnover to job dissatisfaction, contemporary research has expanded its scope to include factors like organizational commitment, climate, and job alternatives.
Aggregate studies suggest that labor market conditions influence overall turnover rates  Terborg (1984). However, when scrutinizing individual decisions to stay or leave an organization, researchers find that general unemployment rates do not exert a direct impact Carsten (1987). Moreover, turnover rates vary across different departments, with industries like accommodation and employee services experiencing higher turnover (around 50%) compared to sectors like software (around 20%) or educational services (over 10%).
Not all turnovers carry the same weight. Initially, involuntary turnover was believed to be within the organization’s control, while uncontrolled turnover presented unique challenges that couldn’t be fully integrated into turnover models. However, managers recognize that not all turnover is detrimental; the departure of underperforming employees may benefit organizational performance, whereas losing high-performing individuals can significantly impact productivity. Hence, it becomes crucial for organizational management to discern between preventable and unpreventable turnover. This distinction allows them to devise targeted strategies to retain valuable talent and mitigate the adverse effects of turnover on organizational effectiveness. Barrick (2005b).
From an economic standpoint, the costs associated with employee turnover are paramount for organizations. Presently, financial statements do not explicitly account for the costs incurred due to voluntary turnover. Instead, these costs are distributed across various processes like recruitment, selection, training, and mentoring. Moreover, the impact of turnover on organizational knowledge continuity is often overlooked. The estimated loss per departing employee can range from thousands of dollars to more than double the employee’s salary, taking into account the loss of organizational knowledge.
Turnover rates also vary across different countries globally. For instance, there are perceptible differences between turnover rates in the United States and European countries. Research findings suggest that Europeans are approximately half as likely to change jobs within a year compared to Americans. In summary, voluntary employee turnover stems from various causes, and maintaining a high employee retention rate is crucial for maintaining a competitive edge over other organizations.
@@
In our research, we have examined various novel factors and their relationships that influence employee turnover, particularly within the IT industry. Given the significant risks associated with IT projects, such as strict deadlines and unclear requirements, project risk factors play a pivotal role in project execution and delivery. These factors can also contribute to job dissatisfaction and subsequent employee turnover.
Furthermore, we have studied the impact of both external/internal social networks on perceived job alternatives, as well as the moderating role of perceived job alternatives between job satisfaction and turnover. We developed a comprehensive questionnaire to assess the influence of these factors within the dynamic IT industry, which was reviewed by project managers and researchers. In our results, we have discussed the influence of these factors mentioned above on employee turnover intentions.
@@ 2 Research Objective
  • To determine the impact of project-specific risks on employee turnover in the context of IT corporations.
  • To determine the mediating role of internal or social links through perceived job alternatives and moderating the relationship in satisfaction of employee job and turnover intentions.
@@ 3 Literature Review
In this section, we define key concepts drawn from existing literature and outline the theories pertaining to employee turnover intentions. Subsequently, we have presented our main ideas that form the basis of the theoretical framework.
@@ 3.1 Theories of Employee Turnover Intentions
a) Theory of Organization Equilibrium March and Simon (1958)
b) Met Expectation Model  Porter (1973)
c) Turnover Process Model  Mobley (1977)
d) Revised Intermediate — Process Model  P W Hom (1984)
e) Structural Model  Price (1981)
f) Expanded Model  Mobley (1979)
g) Multidisciplinary Model  Muchinsky and Morrow (1980)
h) Ferrell and Rusbult: Investment Model  Daniel and Caryl (1981)
@@
J) Cusp Catastrophe Model  John and Michael (1983)
k) Labor Economic Model  Hulin et al. (1985)
l) Unfolding Model Lee (1996)
m) Job embeddedness theory  Mitchell (2001)
@@
In the current competitive environment, attracting and retaining top talent poses a significant challenge for managers. With the expansion of knowledge work, globalization, and technological advancements, the importance of employee retention has heightened compared to previous eras. To mitigate intentional turnover and overall turnover rates, managers implement human resource policies aimed at retaining valuable employees  Fulmer (2003); Hom (2008) Kacmar (2006). Despite the substantial body of research, with over 1500 papers addressing the subject, turnover remains a dynamically evolving field of study. New administrative strategies, changes in job market dynamics, advancements in research methodologies, and technology have all contributed to the evolving nature of turnover research.
@@ 3.2 Time Period Before 1985
There are several major models, proposed before 1985. These models laid the foundation for future studies in employee turnover.
@@
    P W Hom (1984)
One of the seminal theoretical contributions in this field came from  March and Simon (1958). Their groundbreaking work introduced the theory of "organizational equilibrium" in their book published in 1958. They have emphasized the importance of balancing employees, contributions of the organization, and inducements. Central to their theory are two key factors influencing employee harmony: perceived attractiveness and perceived ease of leaving the organization. In contemporary literature, these concepts are commonly referred to as job satisfaction and perceived job alternatives.
@@ 3.3 Turnover Research From 1985 to 1995
From 1985 to 1995, research in the field primarily shifted its focus towards complex organizational and group-level concepts, such as organizational environment, group cohesion, gender composition, and demographics.Unlike the preceding era, which mainly concentrated on individual-level factors, attention expanded to include contextual variables that could have a significant impact on employee turnover.These contextual variables were further categorized into two main groups:
  • Organization macro-level variables
@@
According to  Abelson (1993), organizational culture plays a pivotal role in turnover by shaping employees’ perceptions, leading to withdrawal behaviors. Another study by  Pfeffer (1992) found that pay inequality among university administrators predicts employee turnover, while differences in tenure at the company decrease social integration and influence individual turnover. In addition to macro-level variables like organizational culture and pay systems, researchers also examined other factors such as organizational fit, network centrality, and mentoring.  O’Reilly (1991) concluded that low person-organization fit after 20 months leads to increased turnover. From a social network perspective,  J. Miller McPherson and Drobnic (1992) determined that individuals with more connections within a department are less likely to turnover compared to others.
Another significant advancement during this period was the three-angle conceptualization of organizational commitment by  Meyer (1991). Empirical testing of the three facets of this conceptualization further validated their contributions to understanding turnover-related withdrawal attitudes and behaviors. Researchers also began incorporating recent behaviors into turnover studies, particularly those related to stress, well-being, and ambiguity. Job insecurity and emotional exhaustion were found to be positively linked to turnover intentions.
@@
Overall, this model underscores the dynamics and complexity of the turnover process, suggesting that future research should consider how people leave their jobs. An interesting research contribution is the integrative adaptation and withdrawal model proposed by  Hulin (1991). This model suggests that job dissatisfaction, whether general or specific, triggers a sequence of cognitive and behavioral responses leading to adaptive behaviors. The model argues that withdrawal attitudes are interconnected, and withdrawal behaviors should be viewed as part of a broader pattern encompassing various attitudes, contrasting with previous studies that treated behaviors like turnover and lateness independently.
In conclusion, researchers have broadened their perspective by examining individual-level consequences such as strain and organizational-level consequences like increased turnover among remaining employees at the next job  T. and A. (1990).
@@ 3.4 Turnover Research From 1995 To Present
In the previous decade, there is a considerable theoretical expansion in the area of employee turnover. There are seven major trends of employee turnover in the last decade. All of these trends are described below.
@@
Various theoretical frameworks have contributed to our understanding of employee turnover, enriching the field. However, there remains a gap in comprehending the cooperative aspect of the turnover process. Some studies have delved into individual difference predictors of turnover, exploring both direct and moderating effects. For instance, some researchers have found that personality traits may influence an individual’s decision to leave their job. According to  Barrick (2005b), a person’s confidence and decision-making, when combined with their bio-data, are negatively associated with turnover during the recruitment phase. Additionally,  Barrick (1996) demonstrated that one of the major personality traits, scrupulousness, is negatively correlated with turnover.
@@
Research articles published by  Pelled and Xin (1999) suggest that negative affectivity increases both the intention of leaving and actual turnover. Similarly,  Allen (2005) found that employees with low self-monitoring tendencies and risk reluctance are more likely to turn their desire to leave into actual turnover.
@@
In essential studies aiming to merge different abstract approaches,  Maertz (2004) has connected agreeableness and developed various models. They have demonstrated that their eight turnover factors, previously presented, are systematically associated with four decision categories of turnover. Different groups of job quitters are usually prompted by different forces.
@@
They have identified eight different factors of turnover cognition, which serve as the leading predictors of turnover management. Furthermore, they proposed that these factors intervene in the effects of all other main constructs in the literature. One major finding of their empirical testing is that individuals who quit their jobs without other alternatives have a more negative impact compared to users of other decision types, implying effect-driven, impulsive quitting.
@@
Bauer (2006) published their results on the moderating aspect of extraversion on leader-member interchange and turnover association during new managerial advancement. They found that individuals with low extraversion exhibit a significant negative association between leader-member exchange (LMX) and executive leaving compared to those with high extraversion.
@@
Maertz and their team introduced a single framework that aligns many individual designs to impact turnover, which is quite useful in the research area. With increased importance on organizational change and employee adaptation, rapidly changing environments and stress attitudes have become major areas of behavioral research concerning employee turnover. For example, psychological uncertainty, established as a turnover predictor, was positively linked with turnover intentions and influenced by changing frequency and planning involved in change, as well as transformational change Rafferty and Rafferty (2006). Change acceptance was positively linked with job satisfaction and negatively linked with work displeasure and turnover intentions, aiding in predicting real turnover. Additionally, transformational change positively relates to employee turnover by emphasizing the value retention of employees managed during times of rapid organizational change.
@@
Despite previous researchers including stressors in their models, current research has considered different types of stressors and investigated their benefits. Consistent with previous era of research, hindrance stressors such as role conflict, organizational politics, role overload, and situational constraints are negatively associated with job satisfaction and organizational commitment, while positively related to employee turnover and turnover intentions. On the other hand, challenge stressors like time urgency and task pressure are positively linked to job attitudes and negatively related to turnover intentions  Cavanaugh and Cavanaugh (2000) Podsakoff (2007).
@@
Researchers have also explored stressful events. For instance, Iverson and Pullman found that voluntary turnover and layoffs have different antecedents. Factors which influenced the willingness to leave a job include blue-collar status versus white-collar status (negative correlation), age (negative correlation), intent to leave (positive correlation), and shock (positive correlation). In contrast, layoffs are forecasted by factors such as full-time status versus part-time status (positive correlation), age (positive correlation), and absenteeism (negative correlation). Additionally,  Tepper (2000) stated that there is a positive relation between voluntary turnover and supervisory abuse, mediated by perceptions of organizational justice and moderated by perceived mobility. This relationship is stronger for employees with lower perceived mobility. Furthermore, research by Sims (2005) has shown that sexual harassment has a negative impact on turnover.
@@
Lee (1994) have argued for the need of an alternative theory to explain why and how people leave their companies.They proposed the unfolding model, whose main constituents are shock, script, violations of image, job search, and job satisfaction.Firstly, shock refers to an event that triggers the psychological analysis involved in quitting a job.Aftershock script is a preexisting plan of action required for leaving the job.Image violations occur when personal goals do not align with the organization’s goals and strategies.Additionally, job satisfaction decreases when individuals feel their job does not provide desired technical skills, financial benefits, or knowledge.Finally, job search involves seeking alternative job opportunities and evaluating them.These components of the model evolve over time and integrate to form five unique pathways for individuals to leave their jobs.
@@
In Path 1, shock initiates the execution of a preexisting plan where the individual does not consider alternative job opportunities or their association with the organization. In Path 2, shock prompts the individual to reconsider their association with the organization due to image violations, leading them to leave without considering alternative job opportunities. Path 3 involves shock violating the image of employees, prompting them to compare their current job with other job offers. In Path 4, lower job satisfaction causes the individual to leave the organization without considering alternative job opportunities. Many studies have shown that shock is a significant factor leading individuals to leave their organization. According to  Brooks C. Holtom (2005), shock is the most immediate cause of employee turnover compared to job satisfaction.
@@
Furthermore, a replication of experiments conducted by  Donnelly (2006) extended previous tests on the unfolding model, successfully classifying 86% of samples who left their organization. Their results suggest that most people leave or stay due to economic considerations. Additionally, they found that women experience shock more than men and tend to follow Paths 1, 2, or 3. Research conducted by  Morrell (2004) in Great Britain found that 44% of nurses reported that shocks impacted their decision to turnover.
@@
In summary, there is strong evidence supporting the key concepts of the unfolding model, particularly the concept of shock. However, most of these studies have been conducted on Western data, highlighting the need for similar studies in Eastern contexts.

### v4

@@ Abstract
Employee turnover remains an ongoing issue within high-tech sectors such as IT firms and research centers, where organizational success relies on the skills of their workforce. Intense competition and a scarcity of skilled professionals in the industry contribute to an ongoing demand for highly qualified employees, posing challenges for organizations to retain talent. While numerous studies have explored various factors affecting employee turnover in these industries, their focus often remains on overarching trends rather than specific organizational contexts. In particular, within the software industry, where project-specific risks can affect project success and timely delivery, understanding their influence on job satisfaction and turnover intentions is relevant. This study aims to examine the influence of project risks in the IT industry on job satisfaction and employee turnover intentions. Furthermore, it examines the role of both external and internal social links in shaping perceptions of job alternatives (PJA).
@@
Project risks, Employee turnover intention, Job satisfaction, IT industry
@@ 1 Introduction
Employee turnover presents a notable challenge in the high-tech industry, where skilled personnel are important for success. This is particularly evident in the IT sector, where employees play an important role throughout various project phases, from requirement gathering to development. The departure of main employees can disrupt projects, leading to delays or even failure. While early perspectives attributed turnover to job dissatisfaction, contemporary research has expanded its scope to include factors like organizational commitment, climate, and job alternatives.
Aggregate studies suggest that labor market conditions influence overall turnover rates  75323:2894213. However, when scrutinizing individual decisions to stay or leave an organization, researchers find that general unemployment rates do not exert a direct impact 75323:2894214. Moreover, turnover rates vary across different departments, with industries like accommodation and employee services experiencing higher turnover (around 50%) compared to sectors like software (around 20%) or educational services (over 10%).
Not all turnovers carry the same weight. Initially, involuntary turnover was believed to be within the organization’s control, while uncontrolled turnover presented unique challenges that couldn’t be fully integrated into turnover models. However, managers recognize that not all turnover is detrimental; the departure of underperforming employees may benefit organizational results, whereas losing high-performing individuals can significantly impact productivity. Hence, it becomes relevant for organizational management to discern between preventable and unpreventable turnover. This distinction allows them to devise targeted strategies to retain valuable talent and mitigate the adverse effects of turnover on organizational outcomes. 75323:2894216.
From an economic standpoint, the costs associated with employee turnover are important for organizations. Presently, financial statements do not explicitly account for the costs incurred due to voluntary turnover. Instead, these costs are distributed across various processes like recruitment, selection, training, and mentoring. Moreover, the impact of turnover on organizational knowledge continuity is often overlooked. The estimated loss per departing employee can range from thousands of dollars to more than double the employee’s salary, taking into account the loss of organizational knowledge.
Turnover rates also vary across different countries globally. For instance, there are perceptible differences between turnover rates in the United States and European countries. Research findings suggest that Europeans are approximately half as likely to change jobs within a year compared to Americans. In summary, voluntary employee turnover stems from various causes, and maintaining a high employee retention rate is relevant for maintaining a competitive edge over other organizations.
@@
In our research, we have examined various relatively unexplored factors and their relationships that influence employee turnover, particularly within the IT industry. Given the notable risks associated with IT projects, such as strict deadlines and unclear requirements, project risk factors play a important role in project execution and delivery. These factors can also contribute to job dissatisfaction and subsequent employee turnover.
Furthermore, we have studied the impact of both external/internal social networks on perceived job alternatives (PJA), as well as the moderating role of perceived job alternatives between job satisfaction and turnover. We proposed a structured questionnaire to assess the influence of these factors within the dynamic IT industry, which was reviewed by project managers and researchers. In our results, we have discussed the influence of these factors mentioned above on employee turnover intentions.
@@ 2 Research Objective
  • To examine the impact of project-specific risks on employee turnover in the context of IT corporations.
  • To examine the mediating role of internal or social links through perceived job alternatives and moderating the relationship in satisfaction of employee job and turnover intentions.
@@ 3 Literature Review
In this section, we define main concepts drawn from existing literature and outline the theories pertaining to employee turnover intentions. Subsequently, we have presented our main ideas that form the basis of the theoretical framework.
@@ 3.1 Theories of Employee Turnover Intentions
a) Theory of Organization Equilibrium 75323:2894201
b) Met Expectation Model  75323:2894202
c) Turnover Process Model  75323:2894203
d) Revised Intermediate — Process Model  75323:4296950
e) Structural Model  75323:4296951
f) Expanded Model  75323:2894205
g) Multidisciplinary Model  75323:2894265
h) Ferrell and Rusbult: Investment Model  75323:2894266
@@
J) Cusp Catastrophe Model  75323:2894267
k) Labor Economic Model  75323:2894268
l) Unfolding Model 75323:4302196
m) Job embeddedness theory  75323:2894236
@@
In the current competitive environment, attracting and retaining top talent presents a notable challenge for managers. With the expansion of knowledge work, globalization, and technological advancements, the importance of employee retention has heightened compared to previous eras. To mitigate intentional turnover and overall turnover rates, managers implement human resource policies aimed at retaining valuable employees  75323:2894209; 75323:2894210 75323:2894211. Despite the substantial body of research, with over 1500 papers addressing the subject, turnover remains a dynamically evolving field of study. New administrative strategies, changes in job market dynamics, advancements in research methodologies, and technology have all contributed to the evolving nature of turnover research.
@@ 3.2 Time Period Before 1985
There are several notable models, proposed before 1985. These models laid the foundation for future studies in employee turnover.
@@
    75323:4296950
One of the influential theoretical contributions in this field came from  75323:2894201. Their relevant work introduced the theory of "organizational equilibrium" in their book published in 1958. They have emphasized the importance of balancing employees, contributions of the organization, and inducements. Central to their theory are two main factors influencing employee harmony: perceived attractiveness and perceived ease of leaving the organization. In contemporary literature, these concepts are commonly referred to as job satisfaction and perceived job alternatives.
@@ 3.3 Turnover Research From 1985 to 1995
From 1985 to 1995, research in the field primarily shifted its focus towards complex organizational and group-level concepts, such as organizational environment, group cohesion, gender composition, and demographics.Unlike the preceding era, which mainly concentrated on individual-level factors, attention expanded to include contextual variables that could have a observable impact on employee turnover. These contextual variables were further categorized into two main groups:
  • Organization macro-level variables.
@@
According to  75323:2894217, organizational culture plays a important role in turnover by shaping employees’ perceptions, leading to withdrawal behaviors. Another study by  75323:2894218 found that pay inequality among university administrators predicts employee turnover, while differences in tenure at the company decrease social integration and influence individual turnover. In addition to macro-level variables like organizational culture and pay systems, researchers also examined other factors such as organizational fit, network centrality, and mentoring.  75323:2894219 concluded that low person-organization fit after 20 months leads to increased turnover. From a social network perspective,  75323:2894275 examined that individuals with more connections within a department are less likely to turnover compared to others.
Another notable advancement during this period was the three-angle conceptualization of organizational commitment by  75323:2894220. Empirical testing of the three facets of this conceptualization further assessed their contributions to understanding turnover-related withdrawal attitudes and behaviors. Researchers also began incorporating recent behaviors into turnover studies, particularly those related to stress, well-being, and ambiguity. Job insecurity and emotional exhaustion were found to be positively linked to turnover intentions.
@@
Overall, this model underscores the dynamics and complexity of the turnover process, suggesting that future research should consider how people leave their jobs. An interesting research contribution is the integrative adaptation and withdrawal model proposed by  75323:2894207. This model suggests that job dissatisfaction, whether general or specific, triggers a sequence of cognitive and behavioral responses leading to adaptive behaviors. The model argues that withdrawal attitudes are interconnected, and withdrawal behaviors should be viewed as part of a broader pattern encompassing various attitudes, contrasting with previous studies that treated behaviors like turnover and lateness independently.
In conclusion, researchers have broadened their perspective by examining individual-level consequences such as strain and organizational-level consequences like increased turnover among remaining employees at the next job  75323:2894250.
@@ 3.4 Turnover Research From 1995 To Present
In the previous decade, there is a considerable theoretical expansion in the area of employee turnover. There are seven notable trends of employee turnover in the last decade. All of these trends are described below.
@@
Various theoretical frameworks have contributed to our understanding of employee turnover, enriching the field. However, there remains a gap in comprehending the cooperative aspect of the turnover process. Some studies have delved into individual difference predictors of turnover, exploring both direct and moderating effects. For instance, some researchers have found that personality traits may influence an individual’s decision to leave their job. According to  75323:2894216, a person’s confidence and decision-making, when combined with their bio-data, are negatively associated with turnover during the recruitment phase. Additionally,  75323:2894215 showed that one of the notable personality traits, scrupulousness, is negatively correlated with turnover.
@@
Research articles published by  75323:4301510 suggest that negative affectivity increases both the intention of leaving and actual turnover. Similarly,  75323:2894221 found that employees with low self-monitoring tendencies and risk reluctance are more likely to turn their desire to leave into actual turnover.
@@
In essential studies aiming to merge different abstract approaches,  75323:4301594 has connected agreeableness and proposed various models. They have showed that their eight turnover factors, previously presented, are systematically associated with four decision categories of turnover. Different groups of job quitters are usually prompted by different forces.
@@
They have identified eight different factors of turnover cognition, which serve as the leading predictors of turnover management. Furthermore, they proposed that these factors intervene in the effects of all other main constructs in the literature. One notable finding of their empirical testing is that individuals who quit their jobs without other alternatives have a more negative impact compared to users of other decision types, implying effect-driven, impulsive quitting.
@@
75323:4301595 published their results on the moderating aspect of extraversion on leader-member interchange and turnover association during new managerial advancement. They found that individuals with low extraversion exhibit a notable negative association between leader-member exchange (LMX) and executive leaving compared to those with high extraversion.
@@
Maertz and their team introduced a single framework that aligns many individual designs to impact turnover, which is quite useful in the research area. With increased importance on organizational change and employee adaptation, rapidly changing environments and stress attitudes have become notable areas of behavioral research concerning employee turnover. For example, psychological uncertainty, explored as a turnover predictor, was positively linked with turnover intentions and influenced by changing frequency and planning involved in change, as well as transformational change 75323:2894222. Change acceptance was positively linked with job satisfaction and negatively linked with work displeasure and turnover intentions, aiding in predicting real turnover. Additionally, transformational change positively relates to employee turnover by emphasizing the value retention of employees managed during times of rapid organizational change.
@@
Despite previous researchers including stressors in their models, current research has considered different types of stressors and investigated their benefits. Consistent with previous era of research, hindrance stressors such as role conflict, organizational politics, role overload, and situational constraints are negatively associated with job satisfaction and organizational commitment, while positively related to employee turnover and turnover intentions. On the other hand, challenge stressors like time urgency and task pressure are positively linked to job attitudes and negatively related to turnover intentions  75323:2894223 75323:2894257.
@@
Researchers have also explored stressful events. For instance, Iverson and Pullman found that voluntary turnover and layoffs have different antecedents. Factors which influenced the willingness to leave a job include blue-collar status versus white-collar status (negative correlation), age (negative correlation), intent to leave (positive correlation), and shock (positive correlation). In contrast, layoffs are forecasted by factors such as full-time status versus part-time status (positive correlation), age (positive correlation), and absenteeism (negative correlation). Additionally,  75323:2894224 stated that there is a positive relation between voluntary turnover and supervisory abuse, mediated by perceptions of organizational justice and moderated by perceived mobility. This relationship is stronger for employees with lower perceived mobility. Furthermore, research by 75323:2894225 has shown that harassment has a negative impact on turnover.
@@
75323:2894208 have argued for the need of an alternative theory to explain why and how people leave their companies.They proposed the unfolding model, whose main constituents are shock, script, violations of image, job search, and job satisfaction.Firstly, shock refers to an event that triggers the psychological analysis involved in quitting a job.Aftershock script is a preexisting plan of action required for leaving the job.Image violations occur when personal goals do not align with the organization’s goals and strategies.Additionally, job satisfaction decreases when individuals feel their job does not provide desired technical skills, financial benefits, or knowledge.Finally, job search involves seeking alternative job opportunities and evaluating them.These components of the model evolve over time and integrate to form five unique pathways for individuals to leave their jobs.
@@
In Path 1, shock initiates the execution of a preexisting plan where the individual does not consider alternative job opportunities or their association with the organization. In Path 2, shock prompts the individual to reconsider their association with the organization due to image violations, leading them to leave without considering alternative job opportunities. Path 3 involves shock violating the image of employees, prompting them to compare their current job with other job offers. In Path 4, lower job satisfaction causes the individual to leave the organization without considering alternative job opportunities. Many studies have shown that shock is a notable factor leading individuals to leave their organization. According to  75323:2894277, shock is the most immediate cause of employee turnover compared to job satisfaction.
@@
Furthermore, a replication of experiments conducted by  75323:2894279 extended previous tests on the unfolding model, successfully classifying 86% of samples who left their organization. Their results suggest that most people leave or stay due to economic considerations. Additionally, they found that women experience shock more than men and tend to follow Paths 1, 2, or 3. Research conducted by  75323:2894226 in Great Britain found that 44% of nurses reported that shocks impacted their decision to turnover.
@@
In summary, there is empirical evidence supporting the main concepts of the unfolding model, particularly the concept of shock. However, most of these studies have been conducted on Western data, highlighting the need for similar studies in Eastern contexts.